# Fixation and extinction in time-fluctuating spatially structured metapopulations

### v1

@@ IV.0.2 Strong bottlenecks: $\psi(m,K_{-})<1$
Under these conditions, illustrated in Fig. 6, which depend on $m$ but not on the spatial structure ${\rm G}$, environmental variability generates a succession of strong bottlenecks at a frequency ensuring that the mutant type is the most likely to go extinct in a mean time that is much shorter than the metapopulation mean extinction time. While determining analytically $\Phi$ and $\Theta/\Theta_{E}$ satisfying Eq. (25) is challenging, this can be done efficiently numerically as illustrated by the heatmaps of Fig. 6, and be summarised by maximising the composite quantity $(1-\Phi)\left(1-\Theta/\Theta_{E}\right)$, as shown in Fig. 6(d,h). In the examples of Fig. 6, we find that the near-optimal treatment conditions are $10^{-2}\lesssim\nu\lesssim 1$ and for $K_{-}$ that changes with $m$: $K_{-}\in[2,7]$ for $m=10^{-3}$ and $K_{-}\in[4,9]$ for $m=10^{-4}$. The idealised treatment therefore consists of letting the metapopulation evolve under the near optimal conditions Eq. (25), under which it undergoes a series of strong bottlenecks whose expected outcome is the removal of mutants. Once all mutants are removed, as in Fig. 5(c), the final course of the treatment consists of keeping the metapopulation in the mild environment (with $K=K_{+}$), where $W$ cells would spread and finally take over all the demes. In the example of Fig. 5(c), this would be achieved by setting $K=K_{+}$ after $t\gtrsim 1000$. This idealised treatment, illustrated for clique metapopulations in Fig. 6, qualitatively holds on regular graphs ${\rm G}$, with small influence of the spatial structure on the shape of the heatmap when $m$ is kept fixed, as seen by comparing Figs. 6(e-h) and 10.
@@ V Discussion, generalisations and robustness
Here, we discuss our main results by critically reviewing our assumptions and outline possible generalisations. We have studied the eco-evolutionary dynamics of a metapopulation consisting of $\Omega$ identical demes with the same carrying capacity $K$, containing wild-type $W$ and mutant $M$ cells, that are connected by slow migration and arranged according to regular circulation graphs. While our approach holds for any regular graph, we have specifically considered the examples of cliques (island model), cycles, and square grids (with periodic boundaries) which are all circulation graphs, i.e. the rate of in-flow and out-flow migration is the same at each deme. This has allowed us to consider the effect of both weak and strong local correlations arising respectively in clique and cycle metapopulations. We have analysed the metapopulation dynamics in a static environment where $K$ is constant, and in a time-varying environment where $K$ switches endlessly between $K_{+}$ and $K_{-}<K_{+}$ at a rate $\nu$, see Eq. (5). In static environments, the deme size fluctuates about $K$ and the metapopulation dynamics is characterised by either $M/W$ competition (when $\psi\gg 1$), or by deme extinction (when $\psi<1$). We have used suitable coarse-grained descriptions to analytically characterise the fate of the population in those regimes, see Fig. 2. When, as here, the metapopulation is spatially arranged on circulation graphs, the circulation theorem [41, 47] guarantees that the fixation probability in the competition-dominated regime is independent of the migration rate and the spatial structure. We have also devised a coarse-grained three-state description of the dynamical equilibrium in the crossover regime (where $\psi\gtrsim 1$) where in the long run there is a mixture of occupied demes of type $W$ or $M$ and empty demes, see Sec. F. In time-fluctuating environments, when $K$ switches neither too quickly nor too slowly, each deme is subject to bottlenecks that can be weak when $K_{-}$ is large enough to ensure $\psi(m,K_{-})\gg 1$. Deme extinction can be neglected in the weak bottleneck regime, and we have combined a coarse-grained description with a PDMP approximation to characterise the $W/M$ competition in time-varying environments in the absence of deme extinction. This has allowed us to show that weak bottlenecks lead to a non-monotonic dependence of the mutant fixation probability $\Phi^{{\rm G}}$ and mean fixation time $\Theta^{{\rm G}}$ on the switching rate $\nu$, with an explicit dependence on the migration rate, whereas the spatial structure has an unnoticeable effect on $\Phi^{{\rm G}}$, regardless of spatial correlations, but influences $\Theta^{{\rm G}}$. When demes are subject to strong bottlenecks, metapopulation extinction is a likely outcome under slow and fast switching ($\nu\ll 1$ and $\nu\gg 1$), whereas the overall extinction can be avoided for long periods under intermediate switching, when $W/M$ competition and deme extinction dynamics are coupled. As a hypothetical application, we have considered an idealised treatment for the rapid removal of the mutant conditioned on minimising the risk of overall extinction.
@@
The coarse-grained descriptions adopted in the static and dynamic environments track the dynamics of a single unbreakable cluster of $M$ demes ($M$-cluster). This requires starting from such a cluster, where here we assumed the natural initial condition of a single $M$ deme. Other starting configurations are also possible (e.g. two or more neighbouring $M$ demes) provided that the unbreakable structure of the $M$-cluster is preserved at all times $t\geq 0$. For the sake of concreteness and simplicity, we have focused on a class of regular circulation graphs. As the number of active edges connecting $W$ and $M$ demes in cliques and cycles is known exactly, these graphs are particularly amenable to detailed analysis. In two dimensions, spatial correlations between demes are more complex, and the coarse-grained description of the $M/W$ competition dynamics on a grid has required approximations of the the number of active edges, see Appendix E. A similar approximation is expected to hold on hypercubic lattices (with periodic boundaries). These considerations on the role of the initial condition and spatial structure do not matter when the metapopulation dynamics is dominated by the extinction of demes since these occur randomly. As a consequence, the “idealised treatment” method based on the dynamic coupling of competition and deme extinction to remove a targeted strain is expected to hold on more complex structures, including generic regular graphs.
@@
In this work, we have focused on the biologically relevant regime of slow migration, which is well known to increase the population fragmentation and hence influences its evolution and diversity [61, 88]. Here, the assumption of slow migration is crucial for the coarse-grained description of the metapopulation dynamics, and the values considered in our examples, $m\approx 10^{-5}-10^{-2}$ (see Appendix B), are comparable with those used in microfluidic experimental setups [63]. For $m\gg 1$, the behaviour of a single well-mixed deme is recovered, see Ref. [16]. For intermediate $m$, the dynamics is characterised by coarsening, i.e. the slow growth of domain sizes over time [89, 90]. For the sake of simplicity and without loss of generality, we have assumed that migration occurs without any directional preference and with the same rate for $M$ and $W$. These assumptions can be relaxed and the coarse-grained description be readily generalised to the case of directional and type-specific migration [51], yielding the same qualitative behaviour discussed here for circulation graphs. We note however that asymmetric directional migration significantly affects the evolutionary dynamics on non-circulation graphs, like the star [41, 42, 50, 54]. It would be interesting to study the evolution on these non-circulation graphs in time-varying environments in the case of symmetric and directional migration. For computational tractability, we have chiefly considered metapopulations consisting of $16$ demes of size ranging between $1$ and $200$, which are much smaller systems than most realistic microbial communities. However, we note that with microfluidic devices and single-cell techniques, it is possible to perform spatially structured experiments with $10$ to $100$ cells per microhabitat patch, which are conditions close to those used in our model [63, 91, 92]. Moreover, in-vivo host-associated metapopulations are often fragmented into a limited number of relatively small demes, e.g. $\Omega\approx 25$ and $K\approx 1000$ in mouse lymph nodes [93, 94, 60].
@@ VI Conclusions
Cells evolve in spatially structured settings, where the competition with those nearby is stronger than those further afield, subject to never-ending environmental changes. Spatial structure and environmental variability impact the eco-evolutionary dynamics of microbial populations significantly, but their joint influence is scarcely considered. Mutations frequently arise in cell communities and some may have deleterious effects, e.g. causing the rise of resistance to antibiotics [30, 31, 21, 32]. Motivated by these issues, and inspired by recent advances in microfluidics allowing experiments to track dynamics at the single-cell level [63, 91, 92], we have investigated the prototypical example of a rare mutant having a selective advantage over wild-type resident cells occupying a spatially structured population in time-fluctuating environments. Here, we have considered a class of metapopulation models spatially arranged as regular (circulation) graphs where cells of wild and mutant types compete in a time-fluctuating environment. The metapopulation consists of demes (subpopulations) with the same carrying capacity, connected to each other by slow cell migration. We represent environmental variability by letting the carrying capacity endlessly switch between two values associated to harsh and mild conditions. In this framework, we have studied how migration, spatial structure, and fluctuations influence the probability and time for the mutant or wild-type to take over the metapopulation, and under which conditions extinction of demes and the entire metapopulation occurs. This allows us to identify when environmental variability coupled to demographic fluctuations can be utilised to remove the mutant.
@@
We have first considered the case where demes fluctuate about a constant carrying capacity in static environments. We have thus characterised analytically and using stochastic simulations a regime dominated by the competition between the mutants and wild-type cells, another one where there is deme extinction, as well as a crossover regime combining local competition and extinction. In time-varying environments, various qualitatively different dynamical scenarios arise and environmental fluctuations can significantly influence the evolution of metapopulations. When the rate of switching is neither too slow nor too fast, demes experience bottlenecks and the population is prone to fluctuations or extinction. When the fluctuating carrying capacity remains large and bottlenecks are weak, deme extinction is negligible. The dynamics is thus dominated by the competition between wild-type cells and mutants to invade and take over demes, and eventually the population, which we characterise by devising a suitable coarse-grained description of the individual-based model when migration is slow. This allows us to determine the fixation probability and mean fixation time by combining analytical and computational tools, and to show that these quantities can vary non-monotonically with the environmental switching rate. We find that in the regime of weak bottlenecks, the mutant fixation probability on regular circulation graphs depends on the migration rate, which is in stark contrast with what happens in static environments, while the spatial structure has no noticeable influence. When the carrying capacity is small under harsh conditions, bottlenecks are strong and there is a dynamical coupling of strain competition in the mild environmental state and deme extinction in the harsh environment. This yields rich dynamical scenarios among which we identify a mechanism, expected to hold on any regular graph, driven by environmental variability and fluctuations to efficiently eradicate one strain. As a hypothetical application, we have thus proposed an idealised treatment to remove the mutant, assumed to be deleterious and favoured by selection. We have shown that when each deme is subject to strong bottlenecks at a certain intermediate switching rate the mutant can be efficiently removed by demographic fluctuations arising in the harsh environment without exposing the entire population to a risk of rapid extinction. We have thus determined the near-optimal conditions on the switching rate and bottleneck strength for this idealised treatment and found that these are qualitatively the same on other graphs.
@@
In summary, our analysis sheds further light on the influence of the spatial structure, migration, and fluctuations on the spread of a mutant strain in time-fluctuating environments. We have identified and characterised various dynamical scenarios, displaying a complex dependence on the switching and migration rates. We have also shown how environmental variability and fluctuations can be utilised to achieve desired evolutionary outcomes like the efficient removal of a deleterious mutant. While we have made a number of simplifying assumptions, allowing us to make analytical progress, many of these can be relaxed without affecting the results or the methodology. Our approach holds for arbitrary regular graphs and can be generalised to more complex spatial settings. We therefore believe that the model studied here has numerous potential applications. For instance, it mirrors the in vitro evolution of a mutant across an array of microfluidic devices, where cells migrate between “microhabitat patches” either via microchannels or pipette, with bottlenecks implemented via a strict control of the nutrient level in each device.
@@ A.1 Master equation
As discussed in Sec. II, the individual-based model is a continuous-time multivariate Markov process defined by the reaction and transition rates Eq. (1)-(4). The intra and inter-deme dynamics is encoded in a master equation for the probability $P(\{n_{W},n_{M}\},\xi,t)$ that at time $t$ the metapopulation is in the environmental state $\xi$ and configuration $\{n_{W},n_{M}\}\equiv(\dots,n_{W}(x),n_{M}(x),\dots)$, where $n_{W/M}(x)$ denotes the number of cells of type $W/M$ in deme $x=1,\dots,\Omega$. The master equation (ME) for the metapopulation dynamics subject to environmental switching on a regular graph ${\rm G}=\{{\rm clique,cycle,grid}\}$ of degree (or number of nearest neighbours)
$q_{\rm G}=\begin{cases}\Omega-1,&\text{G$=$~{}clique}\\ 2,&\text{G$=$~{}cycle}\\ 4,&\text{G$=$~{}grid,}\end{cases}$ (26)
reads:
@@
where $y$ n.n. $x$ denotes the sum over the $q_{\rm G}$ neighbours $y$ of the deme $x$ and $P(\dots)=0$ whenever any of $T_{\alpha}^{\pm}$ or $T^{m,{\rm G}}_{\alpha}$ is negative. The shift operators $\mathbb{E}_{\alpha}^{\pm}(x)$ act by raising of decreasing by one the number of cells of type $\alpha$ in deme $x$. For example, $\mathbb{E}_{W}^{\pm}(x)[n_{W}(x)P(\dots,n_{W}(x),n_{M}(x)\dots,\xi,t)]=(n_{W}(x)\pm 1)P(\cdots,n_{W}(x)\pm 1,n_{M}(x)\cdots,\xi,t)$ and $\mathbb{E}_{M}^{\pm}(x)[n_{W}(x)P(\cdots,n_{W}(x),n_{M}(x)\cdots,\xi,t)]=n_{W}(x)P(\cdots,n_{W}(x),n_{M}(x)\pm 1\cdots,\xi,t)$. The first line on the right-hand-side (RHS) of Eq. (27) encodes the inter-deme birth-death dynamics, the second and third lines represent the intra-dynamics via inward and outward migration, and the last line accounts for symmetric random environmental switching. Here, the ME has specifically been formulated in the presence of environmental switching, but its static-environment counterpart is readily obtained from 27: it suffices to set $\nu=0$ and to replace $K(t)$ by a constant carrying capacity $K$, yielding the ME for $P(\{n_{W},n_{M}\},t)$ that is the probability to find the metapopulation in a given state $\{n_{W},n_{M}\}$ at time $t$ (with no environmental dependence). Moreover, by setting $\Omega=1$ and $m=0$ in Eq. (27), the second and third lines on the RHS cancel, we obtain the ME encoding the inter-deme dynamics of a single isolated deme [16, 17].
While the ME Eq. (27) holds for any regular graphs ${\rm G}$, in our examples we consider specifically the regular circulation graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$. The space-dependent individual-based dynamics encoded in the ME Eq. (27) has been simulated using the Monte Carlo method described in Section I. It is worth noting that demographic fluctuations eventually lead to the extinction of the entire metapopulation, in all regimes. However, this phenomenon occurs after a time growing dramatically with the system size, and it can generally not be observed in sufficiently large metapopulations, see Fig. 2(f,g).

### v2

@@ IV.2 Strong bottlenecks: $\psi(m,K_{-})<1$
Under these conditions, illustrated in Fig. 6, which depend on $m$ but not on the spatial structure $\rm G$}, environmental variability generates a succession of strong bottlenecks at a frequency ensuring that the mutant type is the most likely to go extinct in a mean time that is much shorter than the metapopulation mean extinction time. While determining analytically $\Phi$ and $\Theta/\Theta_{E}$ satisfying Eq. (27) is challenging, this can be done efficiently numerically as illustrated by the heatmaps of Fig. 6, and be summarised by maximising the composite quantity $(1-\Phi)\left(1-\Theta/\Theta_{E}\right)$, as shown in Fig. 6(d,h). In the examples of Fig. 6, we find that the near-optimal treatment conditions are $10^{-2}\lesssim\nu\lesssim 1$ and for $K_{-}$ that changes with $m$: $K_{-}\in[2,7]$ for $m=10^{-3}$ and $K_{-}\in[4,9]$ for $m=10^{-4}$. The idealised treatment strategy therefore consists of letting the metapopulation evolve under the near optimal conditions Eq. (27), under which it undergoes a series of strong bottlenecks whose expected outcome is the removal of mutants. Once all mutants are removed, as in Fig. 5(c), the final course of the treatment consists of keeping the metapopulation in the mild environment (with $K=K_{+}$), where $W$ cells would spread and finally take over all the demes. In the example of Fig. 5(c), this would be achieved by setting $K=K_{+}$ after $t\gtrsim 1000$. This idealised treatment strategy, illustrated for clique metapopulations in Fig. 6, qualitatively holds on regular graphs $\rm G$}, with small influence of the spatial structure on the shape of the heatmap when $m$ is kept fixed, as seen by comparing Figs. 6(e-h) and 10.
@@ V Discussion, generalisations and robustness
Here, we discuss our main results by critically reviewing our assumptions and outline possible generalisations. We have studied the eco-evolutionary dynamics of a metapopulation consisting of $\Omega$ identical demes with the same carrying capacity $K$, containing wild-type $W$ and mutant $M$ cells, that are connected by slow migration and arranged according to regular circulation graphs. While our approach holds for any regular graph, we have specifically considered the examples of cliques (island model), cycles, and square grids (with periodic boundaries) which are all circulation graphs, i.e. the rate of in-flow and out-flow migration is the same at each deme. This has allowed us to consider the impact of various graph structures on the metapopulation dynamics. We have analysed the metapopulation dynamics in a static environment where $K$ is constant, and in a time-varying environment where $K$ switches endlessly between $K_{+}$ and $K_{-}<K_{+}$ at a rate $\nu$; see Eq. (6). In static environments, the deme size fluctuates about $K$ and the metapopulation dynamics is characterised by either $M/W$ competition (when $\psi\gg 1$), or by deme extinction (when $\psi<1$). We have used suitable coarse-grained descriptions to analytically characterise the fate of the population in those regimes; see Fig. 2. When, as here, the metapopulation is spatially arranged on circulation graphs, the circulation theorem [41, 47] guarantees that the fixation probability in the competition-dominated regime is independent of the migration rate and the spatial structure. We have also devised a coarse-grained three-state description of the dynamical equilibrium in the crossover regime (where $\psi\gtrsim 1$) where in the long run there is a mixture of occupied demes of type $W$ or $M$ and empty demes; see Appendix F. In time-fluctuating environments, when $K$ switches neither too quickly nor too slowly, each deme is subject to bottlenecks that can be weak when $K_{-}$ is large enough to ensure $\psi(m,K_{-})\gg 1$. Deme extinction can be neglected in the weak bottleneck regime, and we have combined a coarse-grained description with a PDMP approximation to characterise the $W/M$ competition in time-varying environments in the absence of deme extinction. This has allowed us to show that weak bottlenecks lead to a non-monotonic dependence of the mutant fixation probability $\Phi^{{\rm G}}$ and mean fixation time $\Theta^{{\rm G}}$ on the switching rate $\nu$, with an explicit dependence on the migration rate, whereas the spatial structure has an unnoticeable effect on $\Phi^{{\rm G}}$, regardless of spatial correlations, but influences $\Theta^{{\rm G}}$. When demes are subject to strong bottlenecks, metapopulation extinction is a likely outcome under slow and fast switching ($\nu\ll 1$ and $\nu\gg 1$), whereas the overall extinction can be avoided for long periods under intermediate switching, when $W/M$ competition and deme extinction dynamics are coupled. As a hypothetical application, we have considered an idealised treatment strategy for the rapid removal of the mutant conditioned on minimising the risk of overall extinction.
@@
The coarse-grained descriptions adopted in the static and dynamic environments track the dynamics of the number of $M$ demes, which, in the case of the clique and cycle, is a single unbreakable cluster of $M$ demes ($M$-cluster). This requires starting from such a cluster, where here we assumed the natural initial condition of a single $M$ deme. Under these considerations, the number of active edges connecting $W$ and $M$ demes in cliques and cycles is known exactly, making these graphs particularly amenable to detailed analysis. It is also possible to capture the number of active edges exactly for other starting configurations of these graphs (e.g. two or more neighbouring $M$ demes) provided that the unbreakable structure of the $M$-cluster is preserved at all times $t\geq 0$. For the sake of concreteness and simplicity, we have focussed on a class of regular circulation graphs. In two dimensions, spatial correlations between demes are more complex, and the coarse-grained description of the $M/W$ competition dynamics on a grid has required approximations of the the number of active edges; see Appendix E. A similar approximation is expected to hold on hypercubic lattices (with periodic boundaries). These considerations on the role of the initial condition and spatial structure do not matter when the metapopulation dynamics is dominated by the extinction of demes since these occur randomly. As a consequence, the “idealised treatment strategy” based on the dynamic coupling of competition and deme extinction to remove a targeted strain is expected to hold on more complex structures, including generic regular graphs.
@@
In this work, we have focussed on the biologically relevant regime of slow migration, which is well known to increase population fragmentation and hence influences its evolution and diversity [61, 87]. Here, the assumption of slow migration is crucial for the coarse-grained description of the metapopulation dynamics, and the values considered in our examples, $m\approx 10^{-5}-10^{-2}$, are comparable with those used in microfluidic experimental setups [63]. For $m\gg 1$, the behaviour of a single well-mixed deme is recovered; see Ref. [16]. For intermediate $m$, the dynamics is characterised by coarsening, i.e. the slow growth of domain sizes over time [88, 89]. For the sake of simplicity and without loss of generality, we have assumed that migration occurs without any directional preference and with the same rate for $M$ and $W$. These assumptions can be relaxed and the coarse-grained description be readily generalised to the case of directional and type-specific migration [51], yielding the same qualitative behaviour discussed here for circulation graphs. We note however that asymmetric directional migration significantly affects the evolutionary dynamics on non-circulation graphs, like the star [41, 42, 50, 54]. It would be interesting to study the evolution on these non-circulation graphs in time-varying environments in the case of symmetric and directional migration. For computational tractability, we have chiefly considered metapopulations consisting of $16$ demes of size ranging between $1$ and $200$, which are much smaller systems than most realistic microbial communities. However, with microfluidic devices and single-cell techniques, it is possible to perform spatially structured experiments with $10$ to $100$ cells per microhabitat patch, which are conditions close to those used here [63, 90, 91]. Moreover, in-vivo host-associated metapopulations are often fragmented into a limited number of relatively small demes, e.g. $\Omega\approx 25$ and $K\approx 1000$ in mouse lymph nodes [92, 93, 60].
@@ VI Conclusions
Cells evolve in spatially structured settings, where the competition with those nearby is stronger than those further afield, subject to never-ending environmental changes. Spatial structure and environmental variability impact the eco-evolutionary dynamics of microbial populations significantly, but their joint influence is scarcely considered. Mutations frequently arise in cell communities and some may increase cell proliferation while having deleterious effects to a host organism, e.g. leading to cancerous cells or causing resistance to antibiotics [30, 31, 21, 32]. Motivated by these issues, and inspired by recent advances in microfluidics allowing experiments to track dynamics at the single-cell level [63, 90, 91], we have investigated the classic example of a rare mutant having a selective advantage over wild-type resident cells occupying a spatially structured population in time-fluctuating environments. Here, we have considered a class of metapopulation models spatially arranged as regular (circulation) graphs where cells of wild and mutant types compete in a time-fluctuating environment. The metapopulation consists of demes (subpopulations) with the same carrying capacity, connected to each other by slow cell migration. We represent environmental variability by letting the carrying capacity endlessly switch between two values associated to harsh and mild conditions. In this framework, we have studied how migration, spatial structure, and fluctuations influence the probability and time for the mutant or wild-type to take over the metapopulation, and under which conditions extinction of demes and the entire metapopulation occurs. This allows us to identify when environmental variability coupled to demographic fluctuations can be utilised to remove the mutant.
@@
We have first considered the case where demes fluctuate about a constant carrying capacity in static environments. We have thus characterised analytically and using stochastic simulations a regime dominated by the competition between the mutants and wild-type cells, another where there is deme extinction, as well as a crossover regime combining local competition and extinction. In time-varying environments, various qualitatively different dynamical scenarios arise and environmental fluctuations can significantly influence the evolution of metapopulations. When the rate of switching is neither too slow nor too fast, demes experience bottlenecks and the population is prone to fluctuations or extinction. When the fluctuating carrying capacity remains large and bottlenecks are weak, deme extinction is negligible. The dynamics is thus dominated by the competition between wild-type cells and mutants to invade and take over demes, and eventually the population, which we characterise by devising a suitable coarse-grained description of the individual-based model when migration is slow. This allows us to determine the fixation probability and mean fixation time by combining analytical and computational tools, and to show that these quantities can vary non-monotonically with the environmental switching rate. We find that in the regime of weak bottlenecks, the mutant fixation probability on regular circulation graphs depends on the migration rate, which is in stark contrast with what happens in static environments, while the spatial structure has no noticeable influence. When the carrying capacity is small under harsh conditions, bottlenecks are strong and there is a dynamical coupling of strain competition in the mild environmental state and deme extinction in the harsh environment. This yields rich dynamical scenarios among which we identify a mechanism, expected to hold on any regular graph, driven by environmental variability and fluctuations to efficiently eradicate one strain. As a hypothetical application, we have thus proposed an idealised treatment strategy to remove the mutant, assumed to be unwanted and favoured by selection. We have shown that, when each deme is subject to strong bottlenecks at intermediate switching rates, the mutant can be efficiently removed by demographic fluctuations arising in the harsh environment without exposing the entire population to a risk of rapid extinction. We have thus determined the near-optimal conditions on the switching rate and bottleneck strength for this idealised treatment strategy and found that these are qualitatively the same on other graphs.
@@
In summary, our analysis sheds further light on the influence of the spatial structure, migration, and fluctuations on the spread of a mutant strain in time-fluctuating environments. We have identified and characterised various dynamical scenarios, displaying a complex dependence on the switching and migration rates. We have also shown how environmental variability and fluctuations can be utilised to achieve desired evolutionary outcomes like the efficient removal of a pathogenic mutant. While we have made a number of simplifying assumptions, allowing us to make analytical progress, many of these can be relaxed without affecting the results or the methodology. Our approach holds for arbitrary regular graphs and can be generalised to more complex spatial settings. We therefore believe that the model studied here has numerous potential applications. For instance, it mirrors the in vitro evolution of a mutant across an array of microfluidic devices, where cells migrate between “microhabitat patches” either via microchannels or pipette, with bottlenecks implemented via a strict control of the nutrient level in each device.
@@ Acknowledgements.
Table 1: Typical timescales in metapopulation dynamics
Process Timescale for Expression
Deme size growth
Time for deme size to reach
equilibrium
$\ln K$
Deme fixation
Time for fixation
in single deme
$1/s$ [67, 66, 7]
Migration
Expected time between
migration events
$\frac{1}{mK}$
Deme extinction
Mean extinction time
of isolated deme
$\tau_{E}(K)\approx\frac{e^{K}}{K}$
Metapopulation extinction
Extinction time
for full metapopulation
$\theta_{E}\approx\tau_{E}\ln\Omega$
Metapopulation fixation
Unconditional mean
fixation time
Unwieldy expression
Environmental switching
Correlation time of the
dichotomous Markov noise
$\frac{1}{2\nu}$
Table 2: Dynamical regimes and their characteristics
Regime Condition Characteristics Dependence
Static environment
Competition-dominated $\psi\gg 1$
Mutant and wild-type species compete for demes
until fixation of the metapopulation.
$\phi$ independent of $m$ and $\rm{G}$
Extinction-dominated $\psi<1$
Rapid metapopulation extinction
$\theta_{E}$ independent of $\mathrm{G}$
Intermediate $\psi\gtrsim 1$
Competition between demes with stochastic deme
extinction and recolonisation. Fraction of occupied
demes: $\frac{\Omega_{\text{occ}}}{\Omega}\approx 1-\frac{1}{\psi}$
Weak $\mathrm{G}$ dependence of $\phi_{\text{int}}$
Time-varying environment
Weak bottlenecks $\psi(m,K_{-})\gg 1$
Non-monotonic $\nu$-dependence of
$\Phi^{\rm G}$, $\Theta^{\rm G}$
$\Phi^{\rm G}$ depends on $m$
Strong bottlenecks $\psi(m,K_{-})<1$
Extinction likely at $\nu$ extremes.
Mutant removal optimal at intermediate $\nu$
Independent of $\mathrm{G}$
@@ A.1 Master equation
As discussed in Sec. II, the individual-based model is a continuous-time multivariate Markov process defined by the reaction and transition rates Eq. (1)-(4). The intra and inter-deme dynamics is encoded in a master equation for the probability $P(\{n_{W},n_{M}\},\xi,t)$ that at time $t$ the metapopulation is in the environmental state $\xi$ and configuration $\{n_{W},n_{M}\}\equiv(\dots,n_{W}(x),n_{M}(x),\dots)$, where $n_{W/M}(x)$ denotes the number of cells of type $W/M$ in deme $x=1,\dots,\Omega$. The master equation (ME) for the metapopulation dynamics subject to environmental switching on a regular graph ${\rm G}=\{{\rm clique,cycle,grid}\}$ with degrees (or number of nearest neighbours) given by $q_{\text{clique}}=\Omega-1$, $q_{\text{cycle}}=2$, or $q_{\text{grid}}=4$ reads:
@@
where $y$ $n.$ $x$ denotes the sum over the $q_{\rm G}$ neighbours $y$ of the deme $x$ and $P(\dots)=0$ whenever any of $T_{\alpha}^{\pm}$ or $T^{m,{\rm G}}_{\alpha}$ is negative. The shift operators $\mathbb{E}_{\alpha}^{\pm}(x)$ act by raising or decreasing by one the number of cells of type $\alpha$ in deme $x$. For example, $\mathbb{E}_{W}^{\pm}(x)[n_{W}(x)P(\dots,n_{W}(x),n_{M}(x)\dots,\xi,t)]=(n_{W}(x)\pm 1)P(\cdots,n_{W}(x)\pm 1,n_{M}(x)\cdots,\xi,t)$ and $\mathbb{E}_{M}^{\pm}(x)[n_{W}(x)P(\cdots,n_{W}(x),n_{M}(x)\cdots,\xi,t)]=n_{W}(x)P(\cdots,n_{W}(x),n_{M}(x)\pm 1\cdots,\xi,t)$. The first line on the right-hand-side (RHS) of Eq. (28) encodes the intra-deme birth-death dynamics, the second and third lines represent the inter-deme dynamics via inward and outward migration, and the last line accounts for symmetric random environmental switching. Here, the ME has specifically been formulated in the presence of environmental switching, but its static-environment counterpart is readily obtained from Eq. (28): it suffices to set $\nu=0$ and to replace $K(t)$ by a constant carrying capacity $K$, yielding the ME for $P(\{n_{W},n_{M}\},t)$ that is the probability to find the metapopulation in a given state $\{n_{W},n_{M}\}$ at time $t$ (with no environmental dependence). Moreover, by setting $\Omega=1$ and $m=0$ in Eq. (28), the second and third lines on the RHS cancel, we obtain the ME encoding the intra-deme dynamics of a single isolated deme [16, 17].
While the ME Eq. (28) holds for any regular graphs $\rm G$}, in our examples we consider specifically the regular circulation graphs ${\rm G}=\{{\rm clique,cycle,grid}\}$. The space-dependent individual-based dynamics encoded in the ME Eq. (28) has been simulated using the Monte Carlo method described in Appendix I. It is worth noting that demographic fluctuations eventually lead to the extinction of the entire metapopulation, in all regimes. However, this phenomenon occurs after a time growing dramatically with the system size, and it can generally not be observed in sufficiently large metapopulations; see Fig. 2(f).